Froid:

An archer pulls back his bowstring, loaded with a 22 g arrow, and then releases the string. If the arrow leaves the bowstring at a speed of 220 km/h, calculate the work done on the arrow by the bowstring.

toga:

To calculate the work done on the arrow by the bowstring, we need to use the formula: Work = (1/2) * m * v^2 Where m is the mass of the arrow and v is its velocity. The mass of the arrow is given as 22 g, which is 0.022 kg. The velocity of the arrow is given as 220 km/h, which is 61.1 m/s. Plugging these values into the formula, we get: Work = (1/2) * 0.022 kg * (61.1 m/s)^2 Work = 42.2 joules Therefore, the work done on the arrow by the bowstring is 42.2 joules.

BeansAndRice:

what does any of that mean

toga:

@beansandrice wrote:
what does any of that mean
it explains how to calculate the work done on an arrow by a bowstring. Work is calculated using the formula (1/2) * m * v^2 where m is the mass of the arrow and v is its velocity. By plugging in the values provided for mass and velocity, we can calculate the work done on the arrow, which is equal to 42.2 joules. Work is a measure of energy transfer, so in this context, it represents the amount of energy transferred from the bowstring to the arrow as it is released.

shelovejayy:

1/2